# FaithSCAN: Model-Driven Single-Pass Hallucination Detection for Faithful Visual Question Answering

### v1

@@ IV-E3 Uncertainty-aware Attention Fusion
Let $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{R}^{K\times d}$ denote the set of branch-wise evidence embeddings. To aggregate heterogeneous uncertainty sources, FaithSCAN employs a cross-branch attention mechanism:
$\alpha_{k}=\frac{\exp\!\left(\mathbf{w}_{a}^{\top}\tanh\!\left(\mathbf{W}_{a}\mathbf{h}_{k}\right)\right)}{\sum_{j=1}^{K}\exp\!\left(\mathbf{w}_{a}^{\top}\tanh\!\left(\mathbf{W}_{a}\mathbf{h}_{j}\right)\right)},\qquad\mathbf{h}=\sum_{k=1}^{K}\alpha_{k}\mathbf{h}_{k}.$ (20)
To further enhance adaptivity, we introduce a gated residual fusion:
$\mathbf{g}=\sigma\!\left(\mathbf{W}_{g}\mathbf{h}\right),\qquad\widetilde{\mathbf{h}}=\mathbf{h}\odot\mathbf{g}+\mathbf{h},$ (21)
where $\sigma(\cdot)$ denotes the sigmoid function. This formulation enables selective modulation of aggregated uncertainty evidence while preserving the original representation through the residual path.
@@ IV-E4 Hallucination Prediction
where $p\in(0,1)$ denotes the likelihood that the generated answer is hallucinated. FaithSCAN is trained end-to-end using binary cross-entropy loss with model-driven supervision, while the underlying VLM remains frozen.
@@ V-A Datasets
We evaluate our approach on four multimodal datasets with complementary characteristics: HalLoc-VQA [park2025halloc], POPE [li-etal-2023-evaluating], HaloQuest [wang2024haloquest], and VQA v2 [goyal2017vqa], which differ in hallucination types, annotation granularity, and inducing mechanisms. HalLoc-VQA provides token-level supervision for object- and scene-level hallucinations. POPE focuses on object-level hallucinations via direct probing of object existence, HaloQuest emphasizes false-premise and reasoning-induced hallucinations, and VQA v2 serves as a naturalistic evaluation set for real-world visual grounding challenges. Dataset statistics are summarized in Table II; for training and evaluation, we sample 1,000 instances from each split (full HaloQuest test set used) and report averages over three random seeds, ensuring consistent evaluation while controlling computational cost, following common practice in multimodal hallucination and uncertainty analysis [lee2024vhelm, chen-etal-2025-unveiling, park2025glsim, farquhar2024detecting].
@@ V-B Baselines
We compare FaithSCAN with a diverse set of representative baselines for multimodal hallucination detection, focusing on uncertainty-based methods at token, latent, and semantic levels. Token-level baselines include Predictive Entropy (PE) [holtzman2020curious] and Token-level Negative Log-Likelihood (T-NLL) [zhang2023enhancing], measuring per-token predictive uncertainty. Latent-space baselines such as Embedding Variance (EmbVar) [grewal2024improving] and Semantic Embedding Uncertainty (SEU) [grewal2024improving] capture instability in vision-conditioned answer embeddings. Semantic-level baselines include Semantic Entropy (SE) [farquhar2024detecting] and a multimodal adaptation of SelfCheckGPT [manakul2023selfcheckgpt], detecting contradictions across answer pairs. A multimodal $P(\text{True})$ verification baseline [farquhar2024detecting] evaluates answer correctness via image-conditioned yes/no prompts. Finally, a Logistic Regression [hosmer2013applied] classifier trained on our dataset serves as a supervised reference.
@@ V-C Implementation Details and Metrics
Experiments are conducted on three vision–language models: IB-T5-XL, LLaVA-8B, and Qwen3-VL-8B [^1] Full model names: IB-T5-XL (Salesforce/instructblip-flan-t5-xl); LLaVA-8B (llava-hf/llama3-llava-next-8b-hf); Qwen3-VL-8B (Qwen/Qwen3-VL-8B-Instruct). . Uncertainty signals, including token-level log-likelihood, predictive entropy, and hidden representations from a selected transformer layer (see Section VI-C3 for layer-wise analysis), are extracted in a single forward pass during decoding. Hallucination supervision is derived via a model-driven verification scheme using Qwen2.5-VL-32B-Instruct, with labels obtained through low-temperature sampling (temperature 0.1) to validate generated answers. Each uncertainty branch is projected to a 64-dimensional latent space and fused via a lightweight attention module. Models are trained for 20 epochs using AdamW, with learning rate $1\times 10^{-4}$ and batch size 32, employing early stopping on 2×RTX 3090 GPUs, and inference is performed on a single GPU.
Focusing on the I2T setting where evaluation is performed primarily on generated textual answers, we assess hallucination detection using AUROC (ROC) to measure overall ranking quality, AURAC (RAC) to evaluate selective prediction under rejection, F1@Best (F1@B) to report the optimal classification trade-off, and RejAcc@50% (RA@50) to quantify accuracy after rejecting the most uncertain half of predictions [farquhar2024detecting, zhao2025sese, tong2025semantic]. Together, these metrics provide a comprehensive and complementary evaluation of hallucination detection performance from both ranking and decision-making perspectives.
@@
HalLoc-VQA HaloQuest VQA v2 Latency /100 (s)
@@
Logistic 0.662 0.657 0.728 0.666 0.646 0.313 0.858 0.332 0.733 0.740 0.580 0.810 1570.349 0.087
IB-T5-XL FaithSCAN (Ours) 0.778 0.824 0.778 0.860 0.709 0.665 0.684 0.675 0.786 0.813 0.671 0.810 1560.237 0.605
@@
Logistic 0.719 0.771 0.566 0.800 0.792 0.838 0.729 0.857 0.633 0.786 0.432 0.796 1874.895 0.114
LLaVA-8B FaithSCAN (Ours) 0.759 0.801 0.643 0.856 0.681 0.773 0.600 0.762 0.733 0.848 0.536 0.850 1898.950 0.741
@@
Logistic 0.656 0.448 0.493 0.440 0.746 0.294 0.422 0.254 0.610 0.438 0.354 0.416 1689.458 0.102
Qwen3-VL-8B FaithSCAN (Ours) 0.810 0.864 0.671 0.860 0.742 0.886 0.505 0.921 0.738 0.885 0.510 0.920 1741.652 1.191
@@ VI Experiments
In this section, we conduct extensive experiments to evaluate the proposed method from multiple perspectives. Specifically, our experimental study is organized around four research questions. RQ1 (Overall Effectiveness and Generalization): How effective is the proposed method in detecting hallucinations across different vision–language models and datasets covering diverse hallucination types? RQ2 (Model-driven Supervision Analysis): How effective and reliable is the proposed model-driven verification strategy in providing supervision signals for hallucination detection? RQ3 (Internal Representation Analysis): What roles do different internal representations, including token-level uncertainty, visual embeddings, and aligned visual–semantic embeddings, play in hallucination detection, and how do they complement each other? RQ4 (Practical Utility and Interpretability): How does the proposed method perform under practical decision-making scenarios, including its behavior at different recall levels and the interpretability of its internal uncertainty signals?
@@ VI-A Overall Effectiveness and Generalization (RQ1)
We evaluate FaithSCAN under in-distribution (ID) and out-of-distribution (OOD) settings across four datasets, using 1,000 randomly sampled instances per split and averaging results over three runs. FaithSCAN leverages token-level uncertainty signals (log-likelihood, entropy, embeddings) and visual patch features, with sequence-level representations obtained via global average pooling and fused through attention (see Section VI-C for architectural choices).
@@ VI-A1 In-distribution Evaluation
ID results (Table III) show that FaithSCAN consistently outperforms strong baselines, achieving up to 10% AUROC improvement and maintaining top performance across metrics and backbone models. Sampling-based baselines such as SelfCheckGPT and $P(\text{True})$ perform competitively but incur higher inference costs, whereas FaithSCAN achieves strong detection in a single pass. SE underperforms due to long-form, detailed answers that undermine clustering assumptions. Performance further improves with more capable backbones, suggesting richer internal representations provide clearer uncertainty signals.
@@ VI-A2 Out-of-distribution Evaluation
VI-B Model-driven Supervision Analysis (RQ2)
@@
Figure 4: RQ3: AUROC scores for different uncertainty source combinations used in hallucination detection across four datasets and three base models. Labels indicate uncertainty types: T = text-level (log-likelihood, entropy, embedding), P = patch-level visual, and A = cross-modal alignment.
@@ VI-C2 Ablation Study on Token-level Uncertainty Modeling
We evaluate token-level uncertainty signals derived from log-likelihoods, entropies, and hidden embeddings. As shown in Fig. 5, embeddings provide the richest signal, followed by entropy, while log-likelihood contributes marginally. Combining these signals consistently improves performance across datasets and models, with attention-based fusion outperforming simple concatenation. Incorporating CNN-based sequence modeling on embeddings further enhances detection, highlighting that localized anomalies in embeddings capture hallucinations effectively. Based on the results in Fig. 5, we adopt the combination of log-likelihood, entropy, embeddings, global average pooling, and attention as a stable FaithSCAN variant, which is used in all subsequent experiments.
@@ VI-D Practical Utility and Interpretability (RQ4)
To assess interpretability, we analyze token-level features using a gradient-based approach inspired by saliency analysis [simonyan2014deep, shrikumar2017learning]. For input features $\mathbf{x}={x_{t,f}}$ and predicted probability $\hat{y}=\sigma(f_{\theta}(\mathbf{x}))$, attributions are computed as $A_{t,f}=(\partial\hat{y}/\partial x_{t,f})\cdot x_{t,f}$, capturing both a feature’s influence on the prediction and its signal strength, yielding interpretable token-level contributions.
@@
Finally, Fig. 9 provides a qualitative case study using token-level attributions obtained via the grad $\times$ input method. Due to the difficulty of aligning visual features with textual tokens, we focus on text-side feature channels in this analysis. In the illustrated example, Qwen3-VL-8B incorrectly identifies a stool instead of a chair. The key semantic distinction lies in the presence of a backrest, which is overlooked by the model despite rich visual details. FaithSCAN assigns strong positive contributions to the corresponding hallucinated token sequence, accurately capturing the erroneous reasoning process. These highlighted regions align well with human intuition, demonstrating the interpretability and diagnostic value of our approach.

### v2

@@ IV-E3 Uncertainty-aware Attention Fusion
Let $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{R}^{K\times d}$ denote the set of evidence embeddings produced by the evidence encoders. To integrate heterogeneous evidence representations, we adopt a cross-branch attention mechanism:
$\alpha_{k}=\frac{\exp\!\left(\mathbf{w}_{a}^{\top}\mathrm{GELU}\!\left(\mathbf{W}_{a}\mathbf{h}_{k}\right)\right)}{\sum_{j=1}^{K}\exp\!\left(\mathbf{w}_{a}^{\top}\mathrm{GELU}\!\left(\mathbf{W}_{a}\mathbf{h}_{j}\right)\right)},\qquad\mathbf{h}=\sum_{k=1}^{K}\alpha_{k}\mathbf{h}_{k}.$ (20)
This attention computes adaptive weights over branches, allowing the model to emphasize certain evidence representations while aggregating information across all branches.
To further modulate the fused representation, we apply a gated residual transformation:
$\mathbf{g}=\tanh\!\left(\mathbf{W}_{g}\mathbf{h}\right),\qquad\widetilde{\mathbf{h}}=\mathbf{h}\odot\mathbf{g}+\mathbf{h},$ (21)
where the $\tanh(\cdot)$ gate modulates each feature of the fused representation, and the residual connection preserves the original information while allowing adaptive adjustment.
@@ IV-E4 Hallucination Prediction
where $p\in(0,1)$ denotes the likelihood that the generated answer is hallucinated. FaithSCAN is trained end-to-end with model-driven supervision while keeping the underlying VLM frozen, using a binary cross-entropy loss over the hallucination labels.
@@ V-A Datasets
We evaluate our approach on four multimodal datasets with complementary characteristics: HalLoc-VQA [34], POPE [27], HaloQuest [42], and VQA v2 [13], which differ in hallucination types, annotation granularity, and inducing mechanisms. HalLoc-VQA provides token-level supervision for object- and scene-level hallucinations. POPE focuses on object-level hallucinations via direct probing of object existence, HaloQuest emphasizes false-premise and reasoning-induced hallucinations, and VQA v2 serves as a naturalistic evaluation set for real-world visual grounding challenges. Dataset statistics are summarized in Table II; for training and evaluation, we sample 2,000 training instances and 1,000 evaluation instances from each split (full HaloQuest test set used) and report averages over three random seeds, ensuring consistent evaluation while controlling computational cost, following common practice in multimodal hallucination and uncertainty analysis [26, 6, 35, 9].
@@ V-B Baselines
We compare FaithSCAN with a diverse set of representative baselines for multimodal hallucination detection, focusing on uncertainty-based methods at token, latent, and semantic levels. Token-level baselines include Predictive Entropy (PE) [16] and Token-level Negative Log-Likelihood (T-NLL) [46], measuring per-token predictive uncertainty. Latent-space baselines such as Embedding Variance (EmbVar) [14] and Semantic Embedding Uncertainty (SEU) [14] capture instability in vision-conditioned answer embeddings. Semantic-level baselines include Semantic Entropy (SE) [9] and a multimodal adaptation of SelfCheckGPT [33], detecting contradictions across answer pairs. A multimodal $P(\text{True})$ verification baseline [9] evaluates answer correctness via image-conditioned yes/no prompts. Finally, a Logistic Regression [17] classifier trained on our dataset serves as a supervised reference, with implementation details and key hyperparameters provided in Supplementary Material Sec. C.
@@ V-C Implementation Details and Metrics
We evaluate FaithSCAN on three vision–language backbones: IB-T5-XL, LLaVA-8B, and Qwen3-VL-8B [^1] Full model names: IB-T5-XL (Salesforce/instructblip-flan-t5-xl); LLaVA-8B (llava-hf/llama3-llava-next-8b-hf); Qwen3-VL-8B (Qwen/Qwen3-VL-8B-Instruct). . Uncertainty signals, including token-level log-likelihood, predictive entropy, and hidden representations from a selected transformer layer, are extracted in a single forward pass and fused via an attention-based module after projection to 64 dimensions. Hallucination supervision is obtained via model-driven verification using Qwen2.5-VL-32B-Instruct with low-temperature sampling ($T=0.1$). Models are trained with AdamW and binary cross-entropy loss with logits for up to 40 epochs, with early stopping based on validation AUROC. Prompt templates are provided in Supplementary Material Sec. A, and model implementation details are given in Sec. B.
Focusing on the I2T setting where evaluation is performed primarily on generated textual answers, we assess hallucination detection using AUROC (ROC) to measure overall ranking quality, AURAC (RAC) to evaluate selective prediction under rejection, F1@Best (F1@B) to report the optimal classification trade-off, and RejAcc@50% (RA@50) to quantify accuracy after rejecting the most uncertain half of predictions [9, 48, 41]. Together, these metrics provide a comprehensive and complementary evaluation of hallucination detection performance from both ranking and decision-making perspectives.
@@
HalLoc-VQA HaloQuest VQA v2 Latency (s/100 samples)
@@
Logistic 0.669 0.669 0.734 0.674 0.735 0.833 0.855 0.807 0.710 0.760 0.555 0.806 1570.349 0.087
IB-T5-XL FaithSCAN (Ours) 0.754 0.778 0.755 0.792 0.713 0.854 0.870 0.885 0.752 0.820 0.573 0.858 1560.237 0.605
@@
Logistic 0.733 0.788 0.580 0.830 0.834 0.840 0.746 0.857 0.592 0.797 0.384 0.802 1874.895 0.114
LLaVA-8B FaithSCAN (Ours) 0.779 0.823 0.680 0.873 0.827 0.848 0.762 0.902 0.685 0.801 0.494 0.831 1898.950 0.741
@@
Logistic 0.672 0.686 0.496 0.724 0.795 0.887 0.446 0.910 0.675 0.812 0.411 0.810 1689.458 0.102
Qwen3-VL-8B FaithSCAN (Ours) 0.739 0.757 0.569 0.811 0.829 0.941 0.532 0.975 0.695 0.876 0.420 0.884 1741.652 1.191
@@ VI Experiments
In this section, we conduct extensive experiments to evaluate the proposed method from multiple perspectives. Specifically, our experimental study is organized around four research questions. RQ1 (Overall Effectiveness and Generalization): How effective is the proposed method in detecting hallucinations across different vision–language models and datasets covering diverse hallucination types? RQ2 (Quality of Model-driven Supervision Signals): How accurate and dependable are the model-driven supervision signals, and to what extent can they support hallucination detection in a supervised learning setup? RQ3 (Internal Representation Analysis): What roles do different internal representations, including token-level uncertainty, visual embeddings, and aligned visual–semantic embeddings, play in hallucination detection, and how do they complement each other? RQ4 (Practical Utility and Interpretability): How does the proposed method perform under practical decision-making scenarios, and to what extent are its decisions interpretable?
@@ VI-A Overall Effectiveness and Generalization (RQ1)
We evaluate FaithSCAN under in-distribution (ID) and out-of-distribution (OOD) settings across four datasets. FaithSCAN leverages token-level uncertainty signals (log-likelihood, entropy, embeddings) and visual patch features, with sequence-level representations obtained via global average pooling and fused through attention (see Section VI-C for architectural choices).
@@ VI-A1 In-distribution Evaluation
ID results (Table III) show that FaithSCAN reaches or surpasses SOTA performance, achieving up to 8% AUROC improvement and maintaining top performance across metrics and backbone models. Sampling-based baselines such as SEU, SelfCheckGPT and $P(\text{True})$ perform competitively but incur higher sampling and inference costs, whereas FaithSCAN achieves strong detection in a single pass. SE underperforms due to long-form, detailed answers that undermine clustering assumptions. Performance further improves with more capable backbones, suggesting richer internal representations provide clearer uncertainty signals.
@@ VI-A2 Out-of-distribution Evaluation
VI-B Quality of Model-driven Supervision Signals (RQ2)
VI-B1 Human Evaluation
@@
VI-B2 Automated Evaluation
We evaluate the reliability of hallucination labels for answers generated by different VLM backbones using three post-hoc scores: Visual-NLI decisiveness ($s_{\text{nli}}$), stochastic decoding consistency ($s_{\text{stoch}}$), and reflection-based self-consistency ($s_{\text{ref}}$). Their distributions are shown in Fig. 4.
$s_{\text{nli}}$ is high for most samples, indicating that the external model produces stable judgments across NLI rounds, with slightly lower consistency on the more challenging VQA v2. $s_{\text{stoch}}$ shows a similar trend across all three datasets, suggesting the generation model maintains consistent internal confidence across different data distributions. $s_{\text{ref}}$ provides a finer-grained view. Hallucinated samples tend to have lower $s_{\text{ref}}$ because unfaithful outputs make reflection less stable. Differences across datasets also appear: HaloQuest contains some uncertain cases due to false-premise hallucinations, while hallucinations in VQA v2 and HalLoc-VQA are more often clearly judged as entailment or contradiction.
These scores show that model-driven hallucination labels are generally stable, while reflection provides deeper insight into the link between data distributions and hallucination labels, highlighting a promising direction for future study.
Figure 4: Distributions of post-hoc reliability scores on three datasets. Answers are from IB-T5-XL; bars are colored by the original three-way hallucination labels (Sec. IV-C1). Results for other backbones are in Supplementary Material D.2.
Figure 5: RQ3: AUROC scores for different uncertainty source combinations used in hallucination detection across the training sets of four datasets and three base models. Labels indicate uncertainty types: T = text-level (log-likelihood, entropy, embedding), P = patch-level visual, and A = cross-modal alignment.
@@ VI-C2 Ablation Study on Token-level Uncertainty Modeling
We evaluate token-level uncertainty signals derived from log-likelihoods, entropies, and hidden embeddings. As shown in Fig. 6, embeddings provide the richest signal, followed by entropy, while log-likelihood contributes marginally. Combining these signals consistently improves performance across datasets and models, with attention-based fusion outperforming simple concatenation. Incorporating CNN-based sequence modeling on embeddings further enhances detection, highlighting that localized anomalies in embeddings capture hallucinations effectively. Supplementary Material Sec. D.1 presents the ablation study on the attention implementation. Based on these results, we adopt the combination of log-likelihood, entropy, embeddings, global average pooling, and attention as a stable FaithSCAN variant, which is used in all subsequent experiments.
@@ VI-D Practical Utility and Interpretability (RQ4)
To assess interpretability, we analyze token-level features using a gradient-based approach inspired by saliency analysis [39, 38]. For input features $\mathbf{x}={x_{t,f}}$ and predicted probability $\hat{y}=\sigma(f_{\theta}(\mathbf{x}))$, attributions are computed as $A_{t,f}=(\partial\hat{y}/\partial x_{t,f})\cdot x_{t,f}$, capturing both a feature’s influence on the prediction and its signal strength, yielding interpretable token-level contributions.
@@
Finally, Fig. 10 presents a qualitative case study using token-level attributions from the grad $\times$ input method (see Supplementary Material Sec. E). Fused visual features are difficult to map back to the original image in an orderly way, so we focus on text-side feature channels. Potential future analyses could leverage image perturbation techniques, as in [45]. In the example, Qwen3-VL-8B incorrectly identifies a stool instead of a chair, as it overlooks the key distinction of the backrest despite rich visual details. FaithSCAN assigns strong positive contributions to the hallucinated token sequence, accurately capturing the erroneous reasoning process. These highlighted regions align with human intuition, demonstrating the interpretability and diagnostic value of our approach.